When setting up the polymerase chain reaction, approximately 200 ng of DNA should be added in a 50 ul reaction. How many ul of the above (it was 115 ng/ul) undiluted DNA solution should be added to the PCR tube
Answer #1

We need 200 ng DNA in 50 ul of reaction mixture.

115 ng DNA = 1 ul

1 ng DNA = 1/115 ul

200 ng DNA = 1/115 × 200 = 1.739 ul

We need to add 1.739 ul of 115 ng/ul to the reaction mixture.

Please rate high.